(x^2+1)d^2y/dx^2-(2x)dx/dy+2y=6(x^2+1)^2 solve it ODE
1
Expert's answer
2020-12-21T17:20:35-0500

"(x^2+1)y''-2xy'+2y=6(x^2+1)^2"

Set "y=vx"

"\implies y'=v+xv'" ,"y''=2v'+v''" .

Put this into the equation.

"(x^2+1)(2v'+xv'')-2x(v+xv')+2xv=6(x^2+1)^2\\
x(x^2+1)v''+2x^2v'+2v'-2xv-2x^2v'+2xv=6(x^2+1)^2\\
x(x^2+1)v''+2v'=6(x^2+1)^2\\"

Divide through by "x(x^2+1)" .

"v''+ \frac{2}{x(x^2+1)}v'=\frac{6(x^2+1)}{x}\\"

Set "u=v',u'=v''" .

"u'+ \frac{2}{x(x^2+1)}u=\frac{6(x^2+1)}{x}\\"


We will be needing an Integrating factor.


"I.F=e^{\int\frac{2}{x(x^2+1)}dx}=e^{\int\frac{2}{x}dx-{\int\frac{2}{x^2+1}dx}}\\
=e^{2\ln x-\ln x^2+1}=\frac{e^{\ln x}}{e^{\ln x^2+1}}=\frac{x^2}{x^2+1}"


Multiply through by the "I.F"


"\frac{x^2}{x^2+1}\left(u'+ \frac{2}{x(x^2+1)}u\right)=\frac{x^2}{x^2+1}\left(\frac{6(x^2+1)}{x}\\\right)\\
\frac{d(u\frac{x^2}{x^2+1})}{dx}=6x"

Integrate both sides


"u\frac{x^2}{x^2+1}=\int 6xdx\\
u\frac{x^2}{x^2+1}=3x^2+C\\
u=3(x^2+1)+\frac{C(x^2+1)}{x^2}\\
\text{But, } u=v'. \text{So,}"

"v'=3(x^2+1)+\frac{C(x^2+1)}{x^2}\\
\text{Integrate both sides}\\
v=x^3+3x+\frac{C(x^2-1)}{x}+K\\
\text{But, } v=\frac{y}{x}. \text{ So, }\\
y=x^4+3x^2+C(x^2-1)+Kx\\
y=x^4+(3+C)x^2+Kx-C"